I am nowhere close to getting the answer can you please help me with this problem?
I have to solve the system by using the substitution or elimination methods.
x/3-y/5=4
3x/4+ 2y/3=6

Answer by Alan3354(69443)

I have to solve the system by using the substitution or elimination methods.
x/3-y/5=4
3x/4+ 2y/3=6
------------
Get rid of fractions.
x/3 - y/5=4 times 15 --> 5x - 3y = 60
3x/4+ 2y/3=6 times 12 --> 9x + 8y = 72
----------
Get the same coefficient for either x or y. I'll do y.
5x - 3y = 60 times 8 --> 40x - 24y = 480
9x + 8y = 72 times 3 --> 27x + 24y = 216
=============
40x - 24y = 480
27x + 24y = 216
--------------- Add
67x = 696
x = 696/67
----------
Sub for x into any eqn to find y.
5x - 3y = 60
3480/67 - 60 = 3y
y = -180/67